Session: Sep 2000 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (A. Polatidis at Fri Sep 15 10:41:53 METDST 2000) Success
Well, this report is for W313B. The observations were succesful.

Torun (K.Borkowski at Fri Sep 15 11:05:04 METDST 2000) Minor failures
Feedback for W313b: since 19:37 we noticed a "no_1pps" popping up from time to time for the BBC2 channel. The reason is still being investigated (in the next experiment (EC015) the offending note appeared in BBC1 instead).

JIVE (Lorant Sjouwerman at Fri Sep 15 11:16:54 METDST 2000) Success
Sorry, I was not sent an update for this experiment. I just noticed Giovanninis W302c was not on the schedule anymore (Ver 3, 28 Aug) but has been replaced by Prestons W313b (Ver 4, 29 Aug). I hope that this was the only change..

Westerbork (Tony Foley at Fri Sep 15 15:52:50 METDST 2000) Minor failures
Twice we had a problem with formatter synch; once before the start, and once during observing about 20:48. The formatter was resynched at 21:28UT, but check the "gps-fmout" -the effect must have been small. We used the tied array with 14 telescopes.

Effelsberg (A. Kraus at Sat Sep 16 18:04:54 METDST 2000) Minor failures
Some computer problems between UT 0000-0040 probably led to the loss of data.

Noto (C.Trigilio at Wed Sep 20 14:09:12 METDST 2000) Success

Jodrell_Bank (Alastair Gunn at Tue Sep 26 06:02:46 METDST 2000) Success
No known problems.